OpenStudy (anonymous):

verify that the equation is an identity: 1+ sin 2x= (sinx + cosx) ^2

OpenStudy (owlfred):

Hoot! You just asked your first question! Hang tight while I find people to answer it for you. You can thank people who give you good answers by clicking the 'Good Answer' button on the right!

OpenStudy (watchmath):

Expand the rigth hand side into \(\sin^2x+2\sin x\cos x+\cos^2 x\) \((\sin^2 x+\cos^2 x) +2\sin x\cos x\) \(1+\sin(2x)\)

OpenStudy (anonymous):

thank you